I need help solving the missing values
4.3 Here is financial statement information on four not-for-profit clinics: Pittman Rose Beckman Jaffe December 31, 2014 Assets Liabilities Equity $80,000 $100,000 g $150,000 d $75,000 50,000 60,000 45,000 90,000 December 31, 2015: 130,000 180,000 k Assets Liabilities Equity 80,000 e110,000 145,000 500,000 55,000 62,000 h 45,000 During 2015: Total revenues 400,000i Total expenses 330,000 360,000 Fill in the missing values labeled a through l.
Answer #1

Use the following equation for solving the balance sheet related question:

Capital(equity) = Assets - Liabilities

Pittman:

2014

a. Equity = Asset - Liabilities = $80,000 - $50,000 = $30,000

2015

b. Asset = Liabilities + equity = 55000 + 45000 = $100,000

c. Expenses = 330000

Revenue = Expenses + Profit

Increase in Asset in 2015 = 100000 - 80000 = 20000

It is evident that out of $20,000, $5,000 came as a result of increase in liability and and $5,000 came as a result of decrease (sale) of equity.

So the other 10000 increase in assets came from the profit earned during the year.

Therefore Total Revenu = $330,000 + $10,000 = $340,000

Rose:

d. Liabilities (2014)= 100000 - 60000 = $40,000

e. Equity (2015) = 130000 - 62000 = $68,000

f. Total Expenses:

Increase in Assetes = 130000 - 100000 = 30000

increase in Liabilities = 62000 - 40000 = 22,000

Increase in Equity = 68000 - 60000 = 8000

out of 30000 increase in assets, 22000 is funded by increase in liabilities, so 8000 remaining and 8000 increase in equity is from profits (16000) from the year.

Total Expenses = $400,000 - $16,000 = $384,000

Beckmen:

g. Assets(2014) = 75000 + 45000 = $120,000

h. Liabilities (2015) = 180000 - 110000 = $70,000

i. Revenues

Incerease in Assets = 180000 - 120000 = $60,000

Decrease in Liabilities = 70000 - 75000 = -$5,000

Increase in Equity = 110000 - 45000 = $65,000

Increase in asset of 60000 is funded 5000 from liabilities, still 55000 remains unadjusted

Also there is gain in equity of 65000.

These funds would have come from profits from the year = 55000 + 65000 = $120.000

Total Revenues = 360000 + 120000 = $480,000

Jaffe:

j. Liabilities (2014) = 150000 - 90000 = $60,000

k. Assets (2015) = 80000 + 145000 = $225,000

Asset increased from 150000 to 225000 = $75,000

Increase in equity = 145000 - 90000 = $55,000

Increase in Liabilities =80000 - 60000 = $20,000

Increase in Assets is funded 20000 from incrase in liabilities

rest 55000 incrase in asset and 55000 increase in equity is a result of profit during the year = 55000+55000 = $110,000

l. Total Expenses = $500,000 - $110,000 = $390,000